OpenStudy (anonymous):

Determine the equation of the plane in the form z = f(x,y) that is parallel to the vectors <9,6,2> and <−8,−4,−5,>, and passes through the point (5,−5,8) I am unsure where to start. I know I find the cross product, but after that I am lost.

OpenStudy (anonymous):

So the cross product will be your normal to the plane. If you take that normal and dot it with the vector <x-x1,y-y1,z-x1> you should get the equation for the plane.

OpenStudy (anonymous):

Where <x1,y1,z1> is a point in the plane.

OpenStudy (amistre64):

you are given the point; P(x,y,z) and the structure is to find the normal <a,b,c> and fill in the equation: a(x-Px) +b(y-Py) +c(z-Pz) = 0

OpenStudy (anonymous):

Err that should be <x-x1,y-y1,z-z1>. Oh, and the dot product should equal 0.

OpenStudy (amistre64):

<9 , 6 , 2 > x <-8,-4,-5> ------------- < a , b , c >

OpenStudy (anonymous):

a(x-Px) +b(y-Py) +c(z-Pz) <-- is the dot product I mentioned ;p

OpenStudy (amistre64):

:) yes it is

OpenStudy (anonymous):

ok, now i understand that the cross product will be normal to the plane but once i find the components of the cross product, do I subtract the points that it needs to pass through?

OpenStudy (amistre64):

any given vector from a given point is determined by: ( x ,y , z ) -(Px,Py,Pz) ------------ which gives all vecotrs in a plane

OpenStudy (amistre64):

Since any given vector in the plane can be found by a given point P(x,y,z) then by doting all those vectors to the normal gives us the plane itself

OpenStudy (amistre64):

given 2 points for example: (2,3,1) and (1,6,5); the vector between them is: (2,3,1) -(1,6,5) ------- <1,-3,-4>; or < x0-x1 ,y0-y1 ,z0-z1 >

OpenStudy (anonymous):

\[<a,b,c>\cdot<x-x_1, y-y_1, z-z_1> = 0\]\[\implies a(x-x_1)+b(y-y_1)+c(z-z_1)=0\]\[\implies ax + by + cz = ax_1 + by_1 + cz_1\]\[\implies <a,b,c>\cdot<x,y,z> = <a,b,c>\cdot<x_1,y_1,z_1>\] Which that last one is usually how I solve these.

OpenStudy (anonymous):

Less algebra that way. \[<a,b,c>\] is the normal \[<x_1,y_1,z_1>\] is a point in the plane

OpenStudy (amistre64):

all vectors from a point say (2,1,6) on the plane to any other point (x,y,z) are then: ( x ,y ,z ) -( 2 ,1 ,6 ) --------------- <x-2, y-1, z-6>

OpenStudy (anonymous):

^ i see now thanks for your help

OpenStudy (anonymous):

would the equation in terms of z be z = (22x - 29y +159)/12???

OpenStudy (anonymous):

Generally the equation for a plane looks like: ax + by + cz = d

OpenStudy (anonymous):

i know, but i need to express it in terms of z for my hw.

OpenStudy (anonymous):

that's odd, and pretty arbitrary but yes, it'd be (-dax -dby)/c then